OpenStudy (anonymous):

Use the unit circle to determine the value of sec(-12pi)

Nnesha (nnesha):

to make it easy add 2pi into -12pi ntill you get positive value and remember sec =1/cos

OpenStudy (anonymous):

i still dont get it

Nnesha (nnesha):

add 2pi \[\huge\rm -12\pi +2\pi\] full circle is equal to 2pi that's why we have to add 2pi keep adding 2 untill u get positive number

Nnesha (nnesha):

\[-12\pi +2\pi = -10 \] still negative add 2pi again \[-10\pi +2\pi = -8\pi \] add 2pi again

OpenStudy (anonymous):

ok i get 0pi

OpenStudy (anonymous):

so it is not an exact value then

Nnesha (nnesha):

\[-12\pi +2\pi = -10 \] still negative add 2pi again \[-10\pi +2\pi = -8\pi \] add 2pi again \[-8\pi +2\pi = -6\pi \] add 2pi \[-6\pi +2\pi = -4\pi \] \[-4\pi +2\pi = -2\pi \] add 2 pi again and let me know what you get

OpenStudy (anonymous):

i get 0pi

Nnesha (nnesha):

okay 0 is correct

Nnesha (nnesha):

not pi just 0

OpenStudy (anonymous):

so it is undefined

Nnesha (nnesha):

like i told you sec =1/cos look at the unit cricle what is cos at 0

OpenStudy (anonymous):

cos(0)=1

Nnesha (nnesha):

yep right so \[\huge\rm sec =\frac{1}{cos} =sec =\frac{ 1 }{ 1 }\] replace cos by 1 = answer

Nnesha (nnesha):

don't forget the *one* at the numerator !

OpenStudy (anonymous):

if the answer is 0 the is it undefined or not

Nnesha (nnesha):

no

Nnesha (nnesha):

they are asking for value of sec(-12pi) in the unit circle radians is 0 to 2pi(-12pi is not in the unit circle t decrease the value we added 2pi ) degree 0 to 360

Nnesha (nnesha):

so in other words question is find the value fo sec(0)

Nnesha (nnesha):

sec(0) is same as sec(-12pi) both are equal

OpenStudy (anonymous):

no i am just asking in general, does the answer 0 mean undefined

Nnesha (nnesha):

ohh okay no well if the DENOMINATOR is 0 then YES!! answer is undefined

Nnesha (nnesha):

okay here is an example let's say they are asking csc(0) and csc equal to 1/sin and now sin (0) is 0 right

Nnesha (nnesha):

so when you replace sin by 0 your answer would be undefined because there is a zero at the denominator \[\large \rm csc = \frac{ 1}{\sin} = \frac{ 1 }{ 0 }=\rm undefined\]

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ 0 }= undefined ~~~~~~~~~~~~~\frac{ 0 }{ 1}=1\] and just simple 0 doesn't mean undefined